# Harnessing Floquet dynamics for selective metrology in few-qubit systems

### v1

@@ I Introduction
Periodically driven quantum systems, described by Floquet theory, offer a platform to realize novel non-equilibrium phases of matter, most notably discrete time crystals (DTCs), which spontaneously break discrete time-translation symmetry under periodic driving Watanabe and Oshikawa (2015); Zaletel et al. (2023); Sacha and Zakrzewski (2017); Else et al. (2020); Wilczek (2012, 2013); Khemani et al. (2016); von Keyserlingk and Sondhi (2016); Else et al. (2017); Harper et al. (2020); Rovny et al. (2018); Yang and Cai (2021); Yu et al. (2019); Liu et al. (2023); Huang et al. (2018); Vu and Das Sarma (2023); Keßler et al. (2020); Kuroś et al. (2020); Keßler et al. (2021). While such symmetry breaking is forbidden in equilibrium, corroborated by no-go theorems Bruno (2013); Watanabe and Oshikawa (2015), it becomes possible in driven systems when energy absorption is suppressed, e.g., by many-body localization Else et al. (2016) or prethermalization Abanin et al. (2017). It is crucial to emphasize that true time-crystalline order is a many-body phenomenon, requiring a thermodynamic limit to exhibit robust, spontaneous symmetry breaking Yao et al. (2017); Zaletel et al. (2023). Nevertheless, the hallmark signatures of DTCs, specifically a subharmonic response such as period-doubling (PD), can manifest even in minimal, finite-size quantum systems Deshmukh et al. (2025). Experimental realizations of such driven phases have been reported Keßler et al. (2021); Kongkhambut et al. (2022); Wu et al. (2024); Kongkhambut et al. (2024); Skulte et al. (2024), and their dynamics have also been investigated for quantum metrological applications Montenegro et al. (2023); Cabot et al. (2024); Iemini et al. (2024); Yousefjani et al. (2025). Such small systems, accessible on near-term quantum devices, can serve as ideal platforms not for studying true phase transitions, but for harnessing these distinctive dynamical regimes for functional purposes.
@@
Quantum metrology aims to estimate physical parameters with high precision Giovannetti et al. (2011); Sidhu and Kok (2020); Montenegro et al. (2025). In systems influenced by several parameters, the estimation of one parameter can be influenced by fluctuations in others, making careful selection of the sensing regime essential. Periodically driven quantum systems can provide an approach to address this challenge, as their dynamical phases can be tuned to enhance sensitivity to a target parameter while suppressing responses to others. The central thesis of this work is that these dynamical regimes can be exploited to create a highly selective quantum sensor. We demonstrate this principle using a minimal three-qubit Ising model subject to a transverse field, where the dynamical phase serves as a functional switch for metrology. By driving the system into a PD regime, it becomes a powerful parameter filter: its sensitivity to the Ising interaction strength is dramatically enhanced. At the same time, its response to the transverse magnetic field is simultaneously suppressed. This ability to actively select which parameter to sense is a key advantage for quantum metrology in noisy environments.
@@ III.1 Period doubling
where $\hat{\mathrm{U}}_{F}$ denotes the Floquet operator for one driving period. This quantity serves as an order parameter reflecting the collective qubit dynamics of the system. In the PD phase, $\langle\hat{M}_{z}\rangle_{t}$ exhibits oscillations with a period of $2T$, whereas in the trivial phase, such as non-PD, it oscillates with a period $T$ or quickly decays to a steady value, shown in Fig. 2(a).
@@ III.2 PD phase diagram and $\pi$-pairing
As a result, only a fraction of the eigenstates are $\pi$-paired with higher spectral weight, and this fraction varies across the $(h_{x}T,JT)$ parameter space. In summary, while $\pi$-pairing remains the underlying mechanism of PD or approximate PD dynamics in our finite system, the dependence on the initial state, system size, and parameters implies that PD here lacks the robustness to perturbations or scaling that characterizes a genuine DTC. Nevertheless, such finite-size and partially paired regimes still exhibit transient or quasi-periodic PD signatures, providing a minimal model to explore the onset of DTC-like dynamics and their connection to metrological precision.
@@ IV.1 Quantum parameter estimation theory
IV.1.1 Estimation of transverse magnetic field $h_{x}$ and interaction strength $J$
Throughout this work, the QFI is employed as a theoretical benchmark to quantify the ultimate precision limits for estimating the parameters $h_{x}$ and $J$, assuming optimal quantum measurements. Rather than fixing a particular measurement operator, we compute the QFI, which represents the maximum achievable precision attainable over all possible POVMs. The extent to which this bound can be saturated in practice depends on the specific measurement protocol, a point we address in section IV.2.
@@ IV.1.1 Estimation of transverse magnetic field $h_{x}$ and interaction strength $J$
These results highlight that PD dynamics amplify sensitivity to the interaction strength $J$, while the non-PD regime favors estimation of the transverse field $h_{x}$. The PD phase is robust yet consistently poor for $h_{x}$ estimation, offering moderate precision for $J$ without dropping to very low performance, whereas the non-PD phase can yield higher sensitivity but with less predictable outcomes. Therefore, when estimating $h_{x}$, one should carefully choose $J$ to avoid parameter regions exhibiting PD behavior, since the maximum QFI lies in the non-PD regime. This can be understood from the Floquet spectrum: in the PD regime, a quasienergy splitting near a subharmonic bifurcation makes the dynamics highly nonlinear in $J$, enhancing its QFI, while the approximate time-crystalline rigidity suppresses the response to a uniform field $h_{x}$, explaining why $J$ emerges as the more sensitive parameter.
IV.2 Classical Fisher information and choice of observables
@@ V Conclusion
By identifying the boundaries between PD and non-PD regimes in the Floquet phase diagram, one can strategically choose the initial conditions and system parameters to operate within the desired dynamical phase. In particular, testing whether the system lies in the PD regime provides a clear route to achieving high-contrast sensitivity to the interaction strength $J$, accompanied by reduced sensitivity to the transverse field $h_{x}$. Conversely, operation in the non-PD regime does not guarantee such selective filtering: depending on the parameters, it may exhibit comparable or enhanced sensitivity to $h_{x}$. This selective control over metrological response, dictated by the underlying dynamical phase, constitutes the key insight and practical relevance of our work. The significance of this result lies in showing that complex few-qubit dynamics can be harnessed for quantum sensing. This metrological advantage is experimentally accessible through the classical Fisher information of observables such as total magnetization and two-qubit correlations, offering a feasible route to near-term realization.
@@ Acknowledgment
This work is supported by the Scientific and Technological Research Council (TÜBİTAK) of Türkiye under Project Grant No. 123F150. We also acknowledge funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Project No. 435696605 and through the Research Units FOR 5413/1, Grant No. 465199066. This work was also supported by the QuantERA II programme (project CoQuaDis, DFG Grant No. 532763411) which has received funding from the EU H2020 research and innovation programme under GA No. 101017733. We also acknowledge support from the Leverhulme Trust (Grant No. RPG-2024-112). Further support was received by the ERC grant OPEN-2QS (Grant No. 101164443).
@@ Appendix A QFI Scaling for $N=3$
where $a$, $b$, and $c$ are fitting coefficients. We take the second derivative of $F_{Q}$ to find the curvature: $d^{2}F_{Q}/dt^{2}=a$, which quantifies the asymptotic growth of the QFI. Figure 7 shows the time evolution of the QFI for estimation of $h_{x}$ in the PD non-PD regimes. As expected, the non-PD regime exhibits faster QFI growth, indicating higher sensitivity to $h_{x}$ at long times. For the PD regime, we find $a=0.426$, whereas in the non-PD regime $a=5.84$. This indicates that, in this example, the non-PD regime exhibits faster growth of QFI and therefore higher sensitivity to $h_{x}$ at long times. The coefficients $a$, $b$, and $c$ are empirical fit parameters that serve to characterize the temporal behavior of QFI rather than representing universal scalings.
@@ Appendix C QFI curvature for the estimation of $h_{x}$ and $J$ in larger spin systems
For completeness, in Fig. 11 we extend the QFI curvature analysis [cf. Eq. (11)] to larger system sizes $N=4,5,6,7$. The top panels show the curvature of QFI for estimation of $h_{x}$, while the bottom panels correspond to estimation of $J$. These results demonstrate that the trends observed for $N=3$ persist in larger systems. In particular, the estimation of $h_{x}$ is consistently enhanced in the non-PD regime, while the PD regime boosts sensitivity to $J$ in larger system sizes. The PD phase is relatively robust with respect to system size $N$ in the $h_{x}T-JT$ diagram. However, it performs poorly for estimating $h_{x}$, while offering moderate performance for the estimation of $J$. By contrast, the non-PD phase can reach much higher precision for sensing, though its performance is less predictable: it can yield either very good or very poor estimates depending on the parameter selection regime.

### v2

@@ I Introduction
Periodically driven quantum systems, described by Floquet theory, offer a platform to realize novel non-equilibrium phases of matter, most notably discrete time crystals (DTCs), which spontaneously break discrete time-translation symmetry under periodic driving Watanabe and Oshikawa (2015); Zaletel et al. (2023); Sacha and Zakrzewski (2017); Else et al. (2020); Wilczek (2012, 2013); Khemani et al. (2016); von Keyserlingk and Sondhi (2016); Else et al. (2017); Harper et al. (2020); Rovny et al. (2018); Yang and Cai (2021); Yu et al. (2019); Liu et al. (2023); Huang et al. (2018); Vu and Das Sarma (2023); Keßler_2020; Kuroś_2020; Keßler et al. (2021). While such symmetry breaking is forbidden in equilibrium, corroborated by no-go theorems Bruno (2013); Watanabe and Oshikawa (2015), it becomes possible in driven systems when energy absorption is suppressed, e.g., by many-body localization Else et al. (2016) or prethermalization Abanin et al. (2017). It is crucial to emphasize that true time-crystalline order is a many-body phenomenon, requiring a thermodynamic limit to exhibit robust, spontaneous symmetry breaking Yao et al. (2017); Zaletel et al. (2023). Nevertheless, the hallmark signatures of DTCs, specifically a subharmonic response such as period-doubling (PD), can manifest even in minimal, finite-size quantum systems Deshmukh et al. (2025). Experimental realizations of such driven phases have been reported Keßler et al. (2021); Kongkhambut et al. (2022); Wu et al. (2024); Kongkhambut et al. (2024); Skulte et al. (2024), and their dynamics have also been investigated for quantum metrological applications Montenegro et al. (2023); Cabot et al. (2024); Iemini et al. (2024); Yousefjani et al. (2025); Shukla et al. (2025). Such small systems, accessible on near-term quantum devices, can serve as ideal platforms not for studying true phase transitions, but for harnessing these distinctive dynamical regimes for functional purposes.
@@
Quantum metrology aims to estimate physical parameters with high precision Giovannetti et al. (2011); Sidhu and Kok (2020); Montenegro et al. (2025). In systems influenced by several parameters, the estimation of one parameter can be influenced by fluctuations in others, making careful selection of the sensing regime essential. Periodically driven quantum systems can provide an approach to address this challenge, as their dynamical phases can be tuned to enhance sensitivity to a target parameter while suppressing responses to others. The central thesis of this work is that these dynamical regimes can be exploited to create a highly selective quantum sensor. Our analysis in the present work does not rely on the thermodynamic limit but focuses on relevant finite-size systems. In this regime, Floquet dynamics act as an effective filter for parameter sensitivity, enabling selective enhancement or suppression of metrological precision depending on the operating point. We demonstrate this principle using a minimal three-qubit Ising model subject to a transverse field, where the dynamical phase serves as a functional switch for metrology. By driving the system into a PD regime, it becomes a powerful parameter filter: its sensitivity to the Ising interaction strength is dramatically enhanced. At the same time, its response to the transverse magnetic field is simultaneously suppressed. This ability to actively select which parameter to sense is a key advantage for quantum metrology in noisy environments.
@@ III.1 Period doubling
where $\hat{\mathrm{U}}_{F}$ denotes the Floquet operator for one driving period. We focus on the $z$-magnetization, which is aligned with the Ising interaction and the Floquet pulse and therefore directly captures the PD response. This quantity serves as an order parameter reflecting the collective qubit dynamics of the system. In the PD phase, $\langle\hat{M}_{z}\rangle_{t}$ exhibits oscillations with a period of $2T$, whereas in the trivial phase, such as non-PD, it oscillates with a period $T$ or quickly decays to a steady value, shown in Fig. 2(a).
@@ III.2 PD phase diagram and $\pi$-pairing
As a result, only a fraction of the eigenstates are $\pi$-paired with higher spectral weight, and this fraction varies across the $(h_{x}T,JT)$ parameter space. In summary, while $\pi$-pairing remains the underlying mechanism of PD or approximate PD dynamics in our finite system, the dependence on the initial state, system size, and parameters implies that PD here lacks the robustness to perturbations or scaling that characterizes a genuine DTC. While DTCs are defined as robust phases of matter protected against disorder and imperfections, the PD behavior discussed here does not necessarily satisfy all such stability criteria. In particular, we do not claim robustness against disorder or perturbations in this finite-size system. Nevertheless, such finite-size and partially paired regimes still exhibit transient or quasi-periodic PD signatures, providing a minimal model to explore the onset of DTC-like dynamics and their connection to metrological precision. We emphasize that the results presented here pertain to closed, finite-size Floquet unitary dynamics. In this setting, observables can exhibit long-lived oscillations, beats, and revivals arising from discrete quasienergy spacings; therefore, the thermodynamic notion of irreversible Floquet heating toward an infinite-temperature state is not directly addressed. Consequently, the PD dynamics and $\pi$-pairing identified in this work are employed to characterize distinct finite-size dynamical regimes, rather than to establish prethermal lifetimes or quantify heating suppression. A systematic analysis of energy absorption, heating rates, and possible prethermal behavior becomes essential when extending these ideas to larger many-body systems or to experimentally relevant open-system settings. Investigating such effects, including their dependence on the parameters $(h_{x}T,JT)$, constitutes an important direction for future work.
@@ IV.1 Quantum parameter estimation theory
IV.2 Estimation of transverse magnetic field $h_{x}$ and interaction strength $J$
Throughout this work, the QFI is employed as a theoretical benchmark to quantify the ultimate precision limits for estimating the parameters $h_{x}$ and $J$, assuming optimal quantum measurements. Rather than fixing a particular measurement operator, we compute the QFI, which represents the maximum achievable precision attainable over all possible POVMs. The extent to which this bound can be saturated in practice depends on the specific measurement protocol, a point we address in section IV.3.
@@ IV.2 Estimation of transverse magnetic field $h_{x}$ and interaction strength $J$
These results highlight that PD dynamics amplify sensitivity to the interaction strength $J$, while the non-PD regime favors estimation of the transverse field $h_{x}$. The PD phase is robust yet consistently poor for $h_{x}$ estimation, offering moderate precision for $J$ without dropping to very low performance, whereas the non-PD phase can yield higher sensitivity but with less predictable outcomes. Therefore, when estimating $h_{x}$, one should carefully choose $J$ to avoid parameter regions exhibiting PD behavior, since the maximum QFI lies in the non-PD regime. This can be understood from the Floquet spectrum: in the PD regime, a quasienergy splitting near a subharmonic bifurcation makes the dynamics highly nonlinear in $J$, enhancing its QFI, while the approximate time-crystalline rigidity suppresses the response to a uniform field $h_{x}$, explaining why $J$ emerges as the more sensitive parameter. Because our analysis focuses on closed, finite-sized Floquet unitary dynamics, the metrological performance discussed here is governed by the structure of the parameter space in $h_{x}T$ and $hT$, rather than by finite prethermal lifetimes associated with many-body heating.
IV.3 Classical Fisher information and choice of observables
@@ V Conclusion
By identifying the boundaries between PD and non-PD regimes in the Floquet phase diagram, one can strategically choose the initial conditions and system parameters to operate within the desired dynamical phase. In particular, testing whether the system lies in the PD regime provides a clear route to achieving high-contrast sensitivity to the interaction strength $J$, accompanied by reduced sensitivity to the transverse field $h_{x}$. Conversely, operation in the non-PD regime does not guarantee such selective filtering: depending on the parameters, it may exhibit comparable or enhanced sensitivity to $h_{x}$. This selective control over metrological response, dictated by the underlying dynamical phase, constitutes the key insight and practical relevance of our work. The significance of this result lies in showing that complex few-qubit dynamics can be harnessed for quantum sensing. This metrological advantage is experimentally accessible through the CFI of observables such as total magnetization and two-qubit correlations, offering a feasible route to near-term realization.
@@ Acknowledgment
This work is supported by the Scientific and Technological Research Council (TÜBİTAK) of Türkiye under Project Grant No. 123F150. We also acknowledge funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Project No. 435696605 and through the Research Units FOR 5413/1, Grant No. 465199066. This work was also supported by the QuantERA II programme (project CoQuaDis, DFG Grant No. 532763411), which has received funding from the EU H2020 research and innovation programme under GA No. 101017733. We also acknowledge support from the Leverhulme Trust (Grant No. RPG-2024-112). Further support was received by the ERC grant OPEN-2QS (Grant No. 101164443).
DATA AVAILABILITY
The data that support the findings of this article are not publicly available. The data are available from the authors upon reasonable request.
@@ Appendix A QFI Scaling for $N=3$
where $a$, $b$, and $c$ are fitting coefficients. We take the second derivative of $F_{Q}$ to find the curvature: $d^{2}F_{Q}/dt^{2}=a$, which quantifies the asymptotic growth of the QFI. Figure 7 shows the time evolution of the QFI for estimation of $h_{x}$ in the PD and non-PD regimes. As expected, the non-PD regime exhibits faster QFI growth, indicating higher sensitivity to $h_{x}$ at long times. For the PD regime, we find $a=0.426$, whereas in the non-PD regime $a=5.84$. This indicates that, in this example, the non-PD regime exhibits faster growth of QFI and therefore higher sensitivity to $h_{x}$ at long times. The coefficients $a$, $b$, and $c$ are empirical fit parameters that serve to characterize the temporal behavior of QFI rather than representing universal scalings.
@@ Appendix C QFI curvature for the estimation of $h_{x}$ and $J$ in larger spin systems
For completeness, in Fig. 11 we extend the QFI curvature analysis [cf. Eq. (11)] to larger system sizes $N=4,5,6,7$. The top panels show the curvature of QFI for estimation of $h_{x}$, while the bottom panels correspond to estimation of $J$. These results demonstrate that the trends observed for $N=3$ persist in larger systems. In particular, the estimation of $h_{x}$ is consistently enhanced in the non-PD regime, while the PD regime boosts sensitivity to $J$ in larger system sizes. The PD phase is relatively robust with respect to system size $N$ in the $h_{x}T-JT$ diagram. However, it performs poorly for estimating $h_{x}$, while offering moderate performance for the estimation of $J$. By contrast, the non-PD phase can reach much higher precision for sensing, though its performance is less predictable: it can yield either very good or very poor estimates depending on the parameter selection regime. We note that the phase diagrams in Fig. 10 exhibit qualitative differences between odd and even system sizes. This even-odd effect is a genuine finite-size phenomenon and persists irrespective of whether Kac scaling is employed, as discussed in Ref. Ullah et al. (2025). For finite $N$, the parity of the system leads to distinct interference patterns in the Floquet dynamics, which manifest as qualitatively different phase diagrams and curvatures of the QFI, as further illustrated in Fig. 11. Physically, these even-odd differences originate from quantum interference Schleich and Wheeler (1987); Schleich et al. (1988); Lassig and Milburn (1993) and are a well-known feature of finite quantum many-body systems. Similar results have been reported previously in the contexts of quantum metrology Alushi et al. (2025); Ullah et al. (2025) and quantum heat engines Hardal et al. (2018).